A tuning fork vibrates at a frequency of 530 Hz. An out-of-tune piano string vibrates at 535 Hz. How much time separates successive beats?
Answer #1

Beat frequency = f1 - f2 = 5 Hz



Time seperation = 1/delF = 1/5 = 0.2 seconds
Answer #2

f = 1/t

beat fri = 535-530 = 5Hz

so time = 1/5= 0.2sec
Answer #3

Beats=frequency2-frequency2=535-530

Beats=5